OpenStudy (anonymous):

How do you do this? Is writing this as a single logarithm different since its subtraction?

OpenStudy (anonymous):

4logx-6log(x+2)

OpenStudy (anonymous):

first use the power rule

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

alogx = log(x^a) use this on both terms

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

how would you do the 6log(x+2) like that ?

OpenStudy (anonymous):

its simply log ((x+2)^6)

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now simplify the two terms using the rule for subtraction

OpenStudy (anonymous):

I don't know what that is, i just started logarithms yesterday

OpenStudy (anonymous):

see log x - log y = log (x/y) log x + log y= log(xy) can you use these?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

would it be \[\log(\frac{ x^{4} }{ (x+2)^{6} }) \]

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

okay thanks (:

OpenStudy (anonymous):

yw